Sarah is renovating her kitchen and has decided to use tiles on her countertop which is 83cm by 203cm. The tiles are octagonal and square in a pattern where four octagons around one square. The square tiles are 3cm along one side.
(a) Estimate the number of tiles they will need to cover the countertop.
(b) Each square tile costs $0.55 and each octagonal tile costs $0.78. Estimate the cost of the tiles to cover the countertop.

Answer by jorel1380(3719)

The best representation I can make is 9 octagonal tiles to 4 square tiles, as seen here:

click here


The total area for this combination of tiles is: (9 x 2(1+√2)3²) + 4(3²)=79.45 cm². The total area for the countertop is 83 x 203=16849 cm².
16849 / 79.45=212
(a) The number of tiles needed:
212 x 9=1908 octagonal tiles
212 x 4=848 square tiles
(b) 1908 x .78=1488.24
848 x .55=466.4
1488.24+466.4=$1954.64 total cost to cover the countertop. ☺☺☺☺